SOME PROBLEMS OF THE INTERSTELLAR CHEMISTRY OF ORGANIC COMPOUNDS

\underline{WILLIAM KLEMPERER}, Department of Chemistry and Chemical Biology, Harvard University, 12 Oxford Street, Cambridge, MA 02138.

We discuss the problem of “excess” oxygen in molecular clouds with the suggestion that the solid phase is the repository. The nature of heating regions is discussed in terms of the changes in gas phase molecular abundances. We discuss the question of quantum chemistries of diffuse clouds with a view of developing quasistatic models for these interesting well studied regions.